# Capturing Turbulence with Numerical Dissipation: a Simple Dynamical Model for Unresolved Turbulence in Hydrodynamic Simulations

### v2

@@ Abstract
Modeling unresolved turbulence in astrophysical gasdynamic simulations can improve the modeling of other subgrid processes dependent on the turbulent structure of gas: from flame propagation in the interiors of combusting white dwarfs to star formation and chemical reaction rates in the interstellar medium, and non-thermal pressure support of circum- and intergalactic gas. We present a simple method for modeling unresolved turbulence in hydrodynamic simulations via tracking its sourcing by local numerical dissipation and modeling its decay into heat. This method is physically justified by the generic property of turbulent flows that they dissipate kinetic energy at a rate set by the energy cascade rate from large scales, which is independent of fluid viscosity regardless of its nature, be it physical or numerical. We calibrate and test our model against decaying supersonic turbulence simulations. Despite its simplicity, the model quantitatively reproduces multiple non-trivial features of the high-resolution turbulence run: the temporal evolution of the average small-scale turbulence, its dependence on spatial scale, and the slope and scatter of the local correlation between subgrid turbulent velocities, gas densities, and local compression rates. As an example of practical applications, we use our model in isolated galactic disk simulations to model locally variable star formation efficiency at the subresolution scale. In the supersonic, star-forming gas, the new model performs comparably to a more sophisticated model where the turbulent cascade is described by explicit subgrid terms. Our new model is straightforward to implement in many hydrodynamic codes used in galaxy simulations as it utilizes already existing infrastructure to implicitly track the numerical dissipation in such codes.
@@ 1 Introduction
At the same time, many of the relevant small-scale processes affecting global behavior strongly depend on the turbulent structure of gas on unresolved scales. Explicit modeling of unresolved turbulence can substantially improve the modeling of such subgrid processes in a wide range of astrophysical applications and phenomena (see, e.g., miesch15; schmidt-review, for reviews): turbulent flame propagation through white dwarf interior during supernovae type Ia combustion (e.g., niemeyer95; reinecke02), unresolved star formation and feedback in galaxy-scale simulations (braun15; semenov16; kretschmer20), turbulent amplification of magnetic fields in the ISM (liu22), turbulent support of gas during the formation of first black holes (latif13), and non-thermal pressure support of galaxy clusters (scannapieco08-turb; maier09), circumgalactic medium (schmidt21), and intergalactic gas (iapichino11), among others. In this paper, we will use galaxy formation simulations as a practical example. In addition to the above applications, in galaxy simulations, subgrid turbulence models can be also used to account for unresolved gas clumping in cooling, heating, and chemical reaction processes, as well as acceleration and transport of cosmic rays (see, e.g., zweibel17; ruszkowski23, for recent reviews).
@@
The subgrid modeling of turbulence in the above applications is often based on the so-called “Large-Eddy Simulation” (LES) methodology (see sagaut; garnier, for extensive reviews). The general idea of such models relies upon the concept of scale separation where the turbulent flow is split into the large-scale, energy-containing part (large eddies), which is modeled by solving modified hydrodynamic equations, and the small-scale part which is followed with a subgrid model that describes the interactions with the resolved flow. Mathematically, the equations of LES can be obtained by applying a spatial filter to hydrodynamic equations and assuming closure relations for the subgrid terms. Such an approach is widely used in simulations of terrestrial turbulent flows with various applications, from industry and aerospace engineering to urban planning and geophysics simulations.
@@
An alternative approach to such hydrodynamic LES is not to include any explicit subgrid turbulence models at all. Although at first glance, the absence of such subgrid models might appear as a severe limitation of such simulations, the past few decades of practical applications showed that this is not the case: simulations of turbulent flows without explicit models can still produce realistic results on scales sufficiently larger than resolution. This phenomenon has been explained via the concept of Implicit Large-Eddy Simulations (LES), whereby the numerical diffusivity of the chosen hydrodynamic method is argued to implicitly approximate the integral effect of all unresolved scales (e.g., boris92; sytine00; margolin02; grete23; malvadi23, see grinstein for a review).
[…]
@@
To this aim, we propose a hybrid method that combines features of both explicit and implicit LES, which we dub as semi-implicit LES (SLES). In this method, the small-scale turbulent energy is followed explicitly, as is the case with standard LES. However, the turbulence source term is computed implicitly by capturing locally dissipated resolved kinetic energy and converting it into subgrid turbulence. This is motivated by the above feature of turbulent flows that the local dissipation rate is set by the energy cascade rate, which constitutes the basis of ILES. The model is straightforward to implement as it relies on the existing infrastructure of ILES codes, while also making the local estimate of unresolved turbulence readily available as in explicit LES.
The paper is structured as follows. In Section 2, we outline further details of the subgrid turbulence models and describe our implementation. Section 3 presents calibration and tests of the model in a decaying supersonic turbulence setup. Section 4 presents the results of the model applied in a galaxy formation simulation. Section 5 discusses our results, and Section 6 summarizes our conclusions. In addition, Appendix A investigates the convergence of our direct high-resolution simulation of decaying turbulence that we used to calibrate and test our model.
@@ 2 Modeling unresolved turbulence
Figure 1: Schematic overview of different types of subgrid turbulence models: implicit, explicit, and, proposed in this work, semi-implicit Large-Eddy Simulations (LES). The semi-implicit LES model combines the explicit modeling of the subgrid turbulent energy and its decay similar to explicit LES with the implicit source term provided by numerical dissipation akin to implicit LES. The labels in the $x$-axis of the power spectrum cartoon represent driving scale ($L_{\rm drive}$), resolution scale ($\Delta$), and physical dissipation scale ($\eta$). Green colors illustrate the explicit subgrid turbulence energy, $e_{\rm turb}$, while “cascade” and “decay” represent the source and sink terms $\Sigma$ and $\epsilon$ in Equation (1), respectively.
In simulations of turbulent astrophysical flows, it is often infeasible to resolve the physical dissipation scale. Instead, the kinetic energy is dissipated on the scales set by resolution, typically corresponding to $\sim$5–10 cells depending on the accuracy of the solver, with numerical effects kicking in on even larger scales, $\sim$30 cells (e.g., kitsionas09; kritsuk11). The information about the turbulent cascade on smaller scales is lost, as illustrated in the left panel Figure 1. The goal of a subgrid turbulence model is, generally, to predict the properties of the missing small-scale turbulence, indicated in the cartoon with the green color.
Below we briefly review two existing broad types of methods used to model subgrid turbulence, implicit and explicit Large-Eddy Simulations (LES), and introduce our new hybrid method that we dub “semi-implicit LES” (SLES) as it combines the elements from both. Note that all these models rely on the idea of a local kinetic energy cascade continuous in spatial scales. In general, the notion of cascade becomes nuanced for supersonic turbulence due to compressibility effects, but it is still applicable in simulations of isothermal supersonic turbulence, in particular, in the context of galaxy formation and ISM simulations (see Section 5).
As a proof-of-concept for our new model, we have implemented it in the adaptive mesh refinement hydrodynamics and $N$-body code ART (kravtsov99; kravtsov02; rudd08; gnedin11), alongside the previously existing explicit LES model (semenov16). The hydrodynamic fluxes in the ART code are handled by a second-order Godunov-type method (colella-glaz) with a piecewise linear reconstruction of states at the cell interfaces (vanleer79) and a monotonized central slope limiter based on colella85, with a lapidus67 artificial viscosity added in regions with strong density gradients. Adiabatic evolution of subgrid turbulent and thermal energies is followed using the entropy-conserving method described in semenov22, while all extra source and sink terms are added in an operator-split manner. The following subsections provide additional details of our implementation of different subgrid turbulence models.
@@ 2.1 Implicit Large-Eddy Simulation (ILES)
The simplest case is the lack of any explicit model for unresolved scales at all. Only the resolved motions are modeled by solving equations of formally inviscid hydrodynamics and these resolved motions are dissipated into heat at the resolution scale by numerical viscosity. This case is illustrated in the first panel on the right-hand side of Figure 1.
Even though the viscosity in such simulations is purely numerical in nature, it ensures the physically correct dissipation rate equal to the energy cascade rate through the resolved scales. This follows from the generic property of turbulent flows that the dissipation rate does not depend on the magnitude or even the nature of viscosity: the kinetic energy is dissipated by cascading down from large to small scales, while viscosity only sets the smallest scale on which these motions are dissipated locally. Thus, in these simulations, the dissipation occurs on the wrong scale (set by the resolution and solver details) but at the correct physical rate at which that energy would be transferred to unresolved scales if such scales were present. In this sense, the effects of numerical viscosity on the resolved flow can be viewed as the integral effect of unresolved scales, implying that an effective subgrid-scale model is implicitly built into the hydrodynamic method. Such an approach to modeling unresolved turbulence is collectively termed Implicit Large-Eddy Simulations [^1] This term is derived from the class of methods with explicit modeling of subgrid scales, “Large-Eddy Simulations,” which is explained in the next section. (ILES; see grinstein, for an extensive review).
For this concept to work, the hydrodynamic scheme must respect the basic physical properties of the underlying hydrodynamic equations being solved, such as conservation, causality, positivity of positive-definite quantities (e.g., mass and energy), and monotonicity (lack of unphysical oscillations). Conservation, in particular, ensures that all dissipated kinetic energy instantly appears as heat as it bypasses the cascade through all unresolved scales. Therefore, in principle, thermal energy in simulations without explicit subgrid turbulence and no radiative cooling can be interpreted as a combination of thermal and subgrid turbulent motions but practical applications do not make such a distinction. In particular, in simulations of galaxy formation and interstellar medium, this energy is subject to radiative cooling and can be quickly radiated away at the incorrect thermal cooling rate.
@@ 2.2 Explicit Large-Eddy Simulation (LES)
Another approach to modeling unresolved turbulence is by explicitly including the terms describing the interactions between subgrid and resolved scales in the hydrodynamic equations. Such class of methods is called Large-Eddy Simulations and it is particularly widespread in various applications of simulations of terrestrial turbulence such as aerospace engineering, urban planning, and climate modeling, among others (see sagaut; garnier, for reviews). Such explicit models are also used in astrophysical applications (see the Introduction and miesch15 and schmidt-review for reviews), although in, e.g., galaxy formation simulations they are not as well spread (see schmidt14; semenov16; kretschmer20, as examples).
@@
This case is illustrated in the middle of Figure 1. One of the key terms that require a closure relation is the sourcing of subgrid turbulence via the energy cascade from the resolved scales (shown with the blue arrow in the sketch). Such a term acts akin to viscosity: it draws kinetic energy from the resolved motions and transfers it to subgrid turbulence. When this turbulent viscosity is added explicitly, it acts on the flow in parallel with the numerical viscosity, and in principle one needs to make sure that the latter is subdominant, e.g., by selecting a high-order method. In practice, however, if the main goal is to predict the local properties of subgrid turbulence rather than accurately model its backreaction on the flow, such a model can still be used with more diffusive, second-order accurate solvers typical for galaxy formation simulations (e.g., semenov16; kretschmer20). In this case, the local flow quantities that enter closure relations are affected by the resolution, however, this approach is still physically motivated because such resolution effects can be viewed as an implicit model for unresolved scales (see Section 2.1 and kretschmer20).
@@
The first line describes purely adiabatic behavior (advection on the left-hand side and $PdV$ work on the right-hand side) which is directly analogous to thermal energy. The turbulent pressure is related to $e_{\rm turb}$ via an ideal gas-like equation of state, $P_{\rm turb}=2/3\;e_{\rm turb}$, which corresponds to the same adiabatic index as that of the thermal gas, $\gamma=5/3$, because unresolved eddies behave analogously to thermal motions. This non-thermal pressure acts on the gas and enters equations for momentum and total energy together with the thermal pressure. We solve the adiabatic part by using an entropy-conservative scheme described in semenov22
@@
The second line in the equation contains the additional terms that describe the evolution of subgrid turbulence, from left to right: sourcing by cascade from the resolved scales ($\Sigma$), decay into heat ($\epsilon$), turbulence diffusion ($\boldsymbol{\nabla}\boldsymbol{\cdot}\kappa\boldsymbol{\nabla}e_{\rm turb}$), and extra explicit sources such as direct injection by stellar feedback ($S$). In the rest of the section, we discuss the source term $\Sigma$, while the other terms are described in Section 2.4.
@@
The source term describing the energy cascade, $\Sigma$, is the key element of the model as it characterizes the interaction between subgrid and resolved scales. Corresponding terms enter the equations for momentum and total energy as viscosity thereby maintaining the conservative form. To preserve conservation, we apply these terms by modifying momentum and energy fluxes and compute the amount of energy that needs to be added to $e_{\rm turb}$ (i.e., the $\Sigma$ term itself) from the difference between total and kinetic energy arising from these extra fluxes. Note, that the latter difference can be reliably computed only when $e_{\rm turb}$ is not too small compared to $e_{\rm tot}$ and therefore when $e_{\rm turb}<10^{-3}\;e_{\rm tot}$ we instead compute $\Sigma$ explicitly and add it to $e_{\rm turb}$ in an operator-split manner.
@@
In the tests presented below, we use two different closure relations for $\Sigma$. In the supersonic turbulence test (Section 3), we use the non-linear closure proposed and calibrated in supersonic regime by schmidt11; specifically, we use their Equation (8) with $C_{1}=0.02$ and $C_{2}=0.7$. For the galaxy formation tests (Section 4), we use the standard eddy-viscosity closure (specifically, the same equation but with $C_{1}=0.095$ and $C_{2}=0$) because we find that this viscous source term is important in the warm and diffuse transonic ISM, while, in the cold and dense supersonic ISM, subgrid turbulence is mainly controlled by the interplay between adiabatic compression and decay making it only weakly sensitive to the specific choice of the closure (semenov16, see also Section 5 below).
@@
Finally, it is worth pointing out the connection with instantaneous models that estimate subgrid velocity dispersion from the gradients of the resolved velocity field and became popular in recent galaxy formation simulations (e.g., hopkins13-sfrecipe; trebitsch17; marinacci19; vintergatan1). Such a model can be viewed as a stationary limit of LES where time-dependent terms are set to zero (i.e., the left-hand side and diffusion term in Equation) and the level of subgrid turbulence is set by the balance between production and dissipation: $-P_{\rm turb}\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}+\Sigma-\epsilon+S=0$ (see also kretschmer20). We leave a detailed comparison with such models for future work.

### v3

@@ Abstract
Modeling unresolved turbulence in astrophysical gasdynamic simulations can improve the modeling of other subgrid processes dependent on the turbulent structure of gas: from flame propagation in the interiors of combusting white dwarfs to star formation and chemical reaction rates in the interstellar medium, and nonthermal pressure support of circum- and intergalactic gas. We present a simple method for modeling unresolved turbulence in hydrodynamic simulations via tracking its sourcing by local numerical dissipation and modeling its decay into heat. This method is physically justified by the generic property of turbulent flows that they dissipate kinetic energy at a rate set by the energy cascade rate from large scales, which is independent of fluid viscosity, regardless of its nature, be it physical or numerical. We calibrate and test our model against decaying supersonic turbulence simulations. Despite its simplicity, the model quantitatively reproduces multiple nontrivial features of the high-resolution turbulence run: the temporal evolution of the average small-scale turbulence, its dependence on spatial scale, and the slope and scatter of the local correlation between subgrid turbulent velocities, gas densities, and local compression rates. As an example of practical applications, we use our model in isolated galactic disk simulations to model locally variable star formation efficiency at the subresolution scale. In the supersonic, star-forming gas, the new model performs comparably to a more sophisticated model where the turbulent cascade is described by explicit subgrid terms. Our new model is straightforward to implement in many hydrodynamic codes used in galaxy simulations, as it utilizes already existing infrastructure to implicitly track the numerical dissipation in such codes.
@@ 1 Introduction
At the same time, many of the relevant small-scale processes affecting global behavior strongly depend on the turbulent structure of gas on unresolved scales. Explicit modeling of unresolved turbulence can substantially improve the modeling of such subgrid processes in a wide range of astrophysical applications and phenomena (see, e.g., miesch15; schmidt-review, for reviews): turbulent flame propagation through the white dwarf interior during supernova (SN) Type Ia combustion (e.g., niemeyer95; reinecke02), unresolved star formation and feedback in galaxy-scale simulations (braun15; semenov16; kretschmer20), turbulent amplification of magnetic fields in the ISM (liu22), turbulent support of gas during the formation of first black holes (latif13), and nonthermal pressure support of galaxy clusters (scannapieco08-turb; maier09), circumgalactic medium (schmidt21), and intergalactic gas (iapichino11), among others. In this paper, we will use galaxy formation simulations as a practical example. In addition to the above applications, in galaxy simulations, subgrid turbulence models can also be used to account for unresolved gas clumping in cooling, heating, and chemical reaction processes, as well as acceleration and transport of cosmic rays (see, e.g., zweibel17; ruszkowski23, for recent reviews).
@@
The subgrid modeling of turbulence in the above applications is often based on the so-called “Large-Eddy Simulation” (LES) methodology (see sagaut; garnier, for extensive reviews). The general idea of such models relies on the concept of scale separation, where the turbulent flow is split into the large-scale, energy-containing part (large eddies), which is modeled by solving modified hydrodynamic equations, and the small-scale part, which is followed by a subgrid model that describes the interactions with the resolved flow. Mathematically, the equations of LES can be obtained by applying a spatial filter to hydrodynamic equations and assuming closure relations for the subgrid terms. Such an approach is widely used in simulations of terrestrial turbulent flows with various applications, from industry and aerospace engineering to urban planning and geophysics simulations.
@@
An alternative approach to such hydrodynamic LES is not to include any explicit subgrid turbulence models at all. Although at first glance, the absence of such subgrid models might appear as a severe limitation of such simulations, the past few decades of practical applications have shown that this is not the case: simulations of turbulent flows without explicit models can still produce realistic results on scales sufficiently larger than the resolution. This phenomenon has been explained via the concept of Implicit Large-Eddy Simulations (ILES), whereby the numerical diffusivity of the chosen hydrodynamic method is argued to implicitly approximate the integral effect of all unresolved scales (e.g., boris92; sytine00; margolin02; grete23; malvadi23, see grinstein for a review).
The ILES hypothesis is based on the generic property of turbulent flows that the rate of dissipation of kinetic energy is independent of the viscosity of the fluid and instead is set by the rate of energy cascade from large scales. The viscosity controls the spatial scale at which dissipation occurs: at a lower viscosity, the turbulent cascade extends down to smaller scales where viscous terms become important (e.g., kolmogorov41; pope). Thus, as long as the numerical method satisfies generic properties of the underlying hydrodynamic equations—such as conservation, causality, and locality—it will ensure the correct rate of energy transfer from large to small scales, while the flow structure and numerical errors of the scheme will conspire on close-to-resolution scales, ensuring dissipation of all incoming kinetic energy at the correct rate (boris92; grinstein; see also Section 5 for a more detailed discussion). The kinetic energy in such simulations is dissipated at the correct rate but on the wrong scale, which is set by resolution and numerics instead of physical viscosity. Extensive numerical studies show that, in the hydrodynamic codes commonly used in astrophysical applications, numerical dissipation effects kick in on rather large scales, corresponding to $\sim 30$ cells, and become dominant on $\sim 5\text{--}10$ cell scales (e.g., kitsionas09; federrath10; federrath11; kritsuk11; grete23; malvadi23). The existing simulations carried out with such codes without explicit subgrid turbulence models can be interpreted as ILES.
@@
To this aim, we propose a hybrid method that combines features of both explicit and implicit LES, which we dub semi-implicit LES (SLES). In this method, the small-scale turbulent energy is followed explicitly, as is the case with standard LES. However, the turbulence source term is computed implicitly by capturing locally dissipated resolved kinetic energy and converting it into subgrid turbulence. This is motivated by the above feature of turbulent flows that the local dissipation rate is set by the energy cascade rate, which constitutes the basis of ILES. The model is straightforward to implement as it relies on the existing infrastructure of ILES codes, while also making the local estimate of unresolved turbulence readily available, as in explicit LES.
The paper is structured as follows. In Section 2, we outline further details of the subgrid turbulence models and describe our implementation. Section 3 presents the calibration and tests of the model in a decaying supersonic turbulence setup. Section 4 presents the results of the model applied in a galaxy formation simulation. Section 5 discusses our results, and Section 6 summarizes our conclusions. In addition, Appendix A investigates the convergence of our direct high-resolution simulation of decaying turbulence that we used to calibrate and test our model.
@@ 2 Modeling unresolved turbulence
Figure 1: Schematic overview of different types of subgrid turbulence models: implicit (ILES), explicit (LES), and proposed in this work, semi-implicit Large-Eddy Simulations (SLES). The SLES model combines the explicit modeling of the subgrid turbulent energy and its decay, similar to explicit LES, with the implicit source term provided by numerical dissipation, akin to ILES. The labels on the $x$-axis of the power spectrum cartoon represent the driving scale ($L_{\rm drive}$), resolution scale ($\Delta$), and physical dissipation scale ($\eta$). Green colors illustrate the explicit subgrid turbulence energy, $e_{\rm turb}$, while “cascade” and “decay” represent the source and sink terms $\Sigma$ and $\epsilon$ in Equation (1), respectively.
In simulations of turbulent astrophysical flows, it is often infeasible to resolve the physical dissipation scale. Instead, the kinetic energy is dissipated on the scales set by the resolution, typically corresponding to $\sim$5–10 cells depending on the accuracy of the solver, with numerical effects kicking in on even larger scales, $\sim$30 cells (e.g., kitsionas09; kritsuk11). The information about the turbulent cascade on smaller scales is lost, as illustrated in the left panel of Figure 1. The goal of a subgrid turbulence model is, generally, to predict the properties of the missing small-scale turbulence, indicated in the cartoon with the green color.
Below, we briefly review two existing broad types of methods used to model subgrid turbulence, implicit and explicit LES, and introduce our new hybrid method that we dub “semi-implicit LES” (SLES), as it combines the elements from both. Note that all these models rely on the idea of a local kinetic energy cascade continuous in spatial scales. In general, the notion of cascade becomes nuanced for supersonic turbulence due to compressibility effects, but it is still applicable in simulations of isothermal supersonic turbulence, particularly in the context of galaxy formation and ISM simulations (see Section 5).
As a proof of concept for our new model, we have implemented it in the adaptive mesh refinement hydrodynamics and $N$-body code Adaptive Refinement Tree (ART; kravtsov99; kravtsov02; rudd08; gnedin11), alongside the previously existing explicit LES model (semenov16). The hydrodynamic fluxes in the ART code are handled by a second-order Godunov-type method (colella-glaz) with a piecewise linear reconstruction of states at the cell interfaces (vanleer79) and a monotonized central slope limiter based on colella85, with a lapidus67 artificial viscosity added in regions with strong density gradients. Adiabatic evolution of subgrid turbulent and thermal energies is followed using the entropy-conserving method described in semenov22, while all extra source and sink terms are added in an operator-split manner. The following subsections provide additional details of our implementation of different subgrid turbulence models.
@@ 2.1 Implicit Large-Eddy Simulation (ILES)
The simplest case is the lack of any explicit model for unresolved scales at all. Only the resolved motions are modeled by solving equations of formally inviscid hydrodynamics, and these resolved motions are dissipated into heat at the resolution scale by numerical viscosity. This case is illustrated in the first panel on the right side of Figure 1.
Even though the viscosity in such simulations is purely numerical in nature, it ensures the physically correct dissipation rate, equal to the energy cascade rate through the resolved scales. This follows from the generic property of turbulent flows that the dissipation rate does not depend on the magnitude or even the nature of viscosity: the kinetic energy is dissipated by cascading down from large to small scales, while viscosity only sets the smallest scale on which these motions are dissipated locally. Thus, in these simulations, the dissipation occurs on the wrong scale (set by the resolution and solver details) but at the correct physical rate at which that energy would be transferred to unresolved scales if such scales were present. In this sense, the effects of numerical viscosity on the resolved flow can be viewed as the integral effect of unresolved scales, implying that an effective subgrid-scale model is implicitly built into the hydrodynamic method. Such an approach to modeling unresolved turbulence is collectively termed Implicit Large-Eddy Simulations [^1] This term is derived from the class of methods with explicit modeling of subgrid scales, “Large-Eddy Simulations,” which is explained in the next section. (ILES; see grinstein, for an extensive review).
For this concept to work, the hydrodynamic scheme must respect the basic physical properties of the underlying hydrodynamic equations being solved, such as conservation, causality, positivity of positive-definite quantities (e.g., mass and energy), and monotonicity (lack of unphysical oscillations). Conservation, in particular, ensures that all dissipated kinetic energy instantly appears as heat as it bypasses the cascade through all unresolved scales. Therefore, in principle, thermal energy in simulations without explicit subgrid turbulence and no radiative cooling can be interpreted as a combination of thermal and subgrid turbulent motions but practical applications do not make such a distinction. In particular, in simulations of galaxy formation and ISM, this energy is subject to radiative cooling and can be quickly radiated away at the incorrect thermal cooling rate.
@@ 2.2 Explicit Large-Eddy Simulation (LES)
Another approach to modeling unresolved turbulence is by explicitly including the terms describing the interactions between subgrid and resolved scales in the hydrodynamic equations. Such a class of methods is called Large-Eddy Simulations and it is particularly widespread in various applications of simulations of terrestrial turbulence, such as aerospace engineering, urban planning, and climate modeling, among others (see sagaut; garnier, for reviews). Such explicit models are also used in astrophysical applications (see the Introduction and miesch15 and schmidt-review for reviews), although in, e.g., galaxy formation simulations, they are not as widespread (see schmidt14; semenov16; kretschmer20, as examples).
@@
This case is illustrated in the middle of Figure 1. One of the key terms that require a closure relation is the sourcing of subgrid turbulence via the energy cascade from the resolved scales (shown with the blue arrow in the sketch). Such a term acts akin to viscosity: it draws kinetic energy from the resolved motions and transfers it to subgrid turbulence. When this turbulent viscosity is added explicitly, it acts on the flow in parallel with the numerical viscosity, and in principle one needs to make sure that the latter is subdominant, e.g., by selecting a high-order method. In practice, however, if the main goal is to predict the local properties of subgrid turbulence rather than accurately model its backreaction on the flow, such a model can still be used with more diffusive, second-order accurate solvers typical for galaxy formation simulations (e.g., semenov16; kretschmer20). In this case, the local flow quantities that enter closure relations are affected by the resolution; however, this approach is still physically motivated because such resolution effects can be viewed as an implicit model for unresolved scales (see Section 2.1 and kretschmer20).
@@
The first line describes purely adiabatic behavior (advection on the left and $PdV$ work on the right), which is directly analogous to thermal energy. The turbulent pressure is related to $e_{\rm turb}$ via an ideal gas-like equation of state, $P_{\rm turb}=2/3\;e_{\rm turb}$, which corresponds to the same adiabatic index as that of the thermal gas, $\gamma=5/3$, because unresolved eddies behave analogously to thermal motions. This nonthermal pressure acts on the gas and enters equations for momentum and total energy together with the thermal pressure. We solve the adiabatic part by using an entropy-conservative scheme described in semenov22.
@@
The second line in the equation contains the additional terms that describe the evolution of subgrid turbulence: from left to right, sourcing by cascade from the resolved scales ($\Sigma$), decay into heat ($\epsilon$), turbulence diffusion ($\boldsymbol{\nabla}\boldsymbol{\cdot}\kappa\boldsymbol{\nabla}e_{\rm turb}$), and extra explicit sources such as direct injection by stellar feedback ($S$). In the rest of the section, we discuss the source term $\Sigma$, while the other terms are described in Section 2.4.
@@
The source term describing the energy cascade, $\Sigma$, is the key element of the model as it characterizes the interaction between subgrid and resolved scales. Corresponding terms enter the equations for momentum and total energy as viscosity, thereby maintaining the conservative form. To preserve conservation, we apply these terms by modifying momentum and energy fluxes and compute the amount of energy that needs to be added to $e_{\rm turb}$ (i.e., the $\Sigma$ term itself) from the difference between total and kinetic energy arising from these extra fluxes. Note that the latter difference can be reliably computed only when $e_{\rm turb}$ is not too small compared to $e_{\rm tot}$ and therefore, when $e_{\rm turb}<10^{-3}\;e_{\rm tot}$, we instead compute $\Sigma$ explicitly and add it to $e_{\rm turb}$ in an operator-split manner.
@@
In the tests presented below, we use two different closure relations for $\Sigma$. In the supersonic turbulence test (Section 3), we use the nonlinear closure proposed and calibrated in the supersonic regime by schmidt11; specifically, we use their Equation (8) with $C_{1}=0.02$ and $C_{2}=0.7$. For the galaxy formation tests (Section 4), we use the standard eddy-viscosity closure (specifically, the same equation but with $C_{1}=0.095$ and $C_{2}=0$) because we find that this viscous source term is important in the warm and diffuse transonic ISM, while, in the cold and dense supersonic ISM, subgrid turbulence is mainly controlled by the interplay between adiabatic compression and decay making it only weakly sensitive to the specific choice of the closure (semenov16, see also Section 5 below).
@@
Finally, it is worthwhile to point out the connection with instantaneous models that estimate subgrid velocity dispersion from the gradients of the resolved velocity field and have become popular in recent galaxy formation simulations (e.g., hopkins13-sfrecipe; trebitsch17; marinacci19; vintergatan1). Such a model can be viewed as a stationary limit of LES where time-dependent terms are set to zero (i.e., the left-hand side and diffusion term in Equation (1)) and the level of subgrid turbulence is set by the balance between production and dissipation: $P_{\rm turb}\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}+\Sigma-\epsilon+S=0$ (see also kretschmer20). We leave a detailed comparison with such models for future work.